# Enhancing Reusability of Learned Skills for Robot Manipulation via Gaze and Bottleneck

### v2

@@ III-C1 Gaze Prediction Model
For gaze-based object manipulation, it is necessary for the robot to acquire gaze information during both training and inference. In response, Kim et al. proposed a method that takes advantage of the common practice of collecting real-robot demonstrations via human teleoperation [12]. During demonstration collection, the teleoperator’s gaze is measured simultaneously, and this gaze data is then used as supervision for a gaze prediction model, which predicts gaze positions across the entire image. The gaze positions estimated by the gaze prediction model are used during both training and inference.
We adopt a similar approach in this study. In particular, we employ an architecture similar to the one used in DAA [15]. As shown in the left side of Figure 5, we replace the image token extraction via EfficientNetV2 with DINOv2 [17], a newer and more powerful feature extraction model, and replace the cross-attention layers (originally intended for language inputs) with a set of four-layer multilayer perceptrons (MLPs) that map each of the 768-dimensional tokens to a single probability value. To intentionally transition the gaze to the next gaze landmark upon completing each sub-task, we prepare one such MLP for each sub-task and switch these MLPs based on the sub-task index $i_{\mathrm{seg}}$. At inference time, the sub-task index $i_{\mathrm{seg}}$ is initialized to zero at the beginning and incremented based on the progress $c_{t}$ output by the action policy model, indicating the completion of each sub-task. By managing gaze transitions using $c_{t}$ estimated from the gaze-centered point cloud, which solely captures the relative spatial relationships between the object and end-effector (or grasped object), the robot can perform gaze transitions at the correct timing even under unseen conditions.
@@ III-C2 Action Policy Model (i)
The pixel-space gaze position predicted by the gaze prediction model is transformed into 3D coordinates using stereo-based depth estimation. From the stereo-derived point cloud, we then crop a cubic region (20 cm on each side in this study) centered on the 3D gaze coordinates, which is referred to as the gaze-centered point cloud. This gaze-centered point cloud is embedded as a sequence of tokens using a PointTransformer [28] and then passed to a Transformer encoder (Figure 5 right). The Transformer encoder includes a CLS token (classification token, similar to ViT [6]) that aggregates information from all point-cloud tokens into a feature vector $f_{\mathrm{pcd}}\in\mathbb{R}^{512}$. All output tokens, including the CLS token, are then fed into a Transformer decoder, which generates both the action sequence $a_{t:t+H}\in\mathbb{R}^{H\times 14}$ for the left and right end-effector over $H$ future time steps and the progress $c_{t}\in\mathbb{R}^{N_{seg}}$ for controlling gaze transitions in each sub-task. Here, each end-effector has 7 degrees of freedom including gripper angle, and $N_{seg}$ denotes the number of sub-tasks. As the input vision is gaze-centered, the output actions are represented as relative end-effector poses, with the current end-effector pose serving as the origin of the coordinate frame. By computing these actions in relative coordinates, the learned dexterous actions remain reusable even when the absolute position of the object changes.
@@ III-C3 Action Policy Model (ii)
After predicting the bottleneck pose, the corresponding reaching trajectory is generated. In this context, a simple straight-line movement sometimes suffices for the reaching motion, because the movement from the current end-effector pose to the bottleneck pose is non-contact and relatively coarse. However, for some tasks the reaching trajectory consistently takes a particular shape. For example, when placing an object into a deep container, the reaching motion toward the container consistently forms an upwardly convex curve. To capture such a characteristic shape without introducing unnecessary complexity that could compromise the model’s generalization, we generate the reaching trajectory from the current end-effector pose to the bottleneck pose by using a first-order Bézier curve. Furthermore, because the endpoint of the Bézier curve is fixed at the bottleneck pose, the robot can accurately reach the bottleneck without incurring a compounding error [29] that accumulates over time.
As illustrated in the lower-right portion of Figure 5, we predict a 7-dimensional bezier vector that determines the shape of the first-order Bézier curve for each end-effector. A first-order Bézier curve is defined by its two endpoints and a single control point. Because the curve lies in a 7-dimensional end-effector pose space, the combined control points for both arms are in 14 dimensions. Rather than predicting the control point directly, we first estimate the displacement—bezier vector—from the mean pose of the start (current end-effector pose) and end (bottleneck pose) poses, and then add this bezier vector to the mean pose to obtain the control point (Figure 5 bottom-right). For the training, we approximate each reaching motion in the demonstrations by fitting them with first-order Bézier curves, and then use the resulting bezier vectors from these curves for supervision.
@@ IV-A Robot System
We used a dual-arm robot system designed for imitation learning via human teleoperation. In this system, a human operator remotely controls the robot while observing its surrounding environment through a head-mounted display (HMD). During this operation, the operator’s gaze data were recorded in sync with the video feed displayed on the HMD. The system is compatible with both a physical dual-arm robot—consisting of two UR5 (Universal Robots Inc.) arms—and its simulated counterpart, allowing the collection of teleoperated demonstration data in both real and virtual environments. The recorded gaze data were output as pixel coordinates corresponding to the images displayed on the HMD. The same single stereo camera (ZED Mini, Stereolabs Inc.) was used for both the teleoperation and inference phases, and we directly utilized the depth images generated by the deep learning–based depth estimation algorithm provided by ZED SDK (Stereolabs Inc.). The time-series demonstration data used for training were recorded at 10 Hz.
@@ IV-B Task Setup
In this study, we conducted imitation learning experiments on three tasks (two real and one simulated) to evaluate the reusability of skills acquired by GazeBot in comparison with conventional models. To quantitatively assess the generalization performance, we collected demonstrations under controlled conditions by restricting the object positions and initial end-effector poses within predefined regions, thus clearly defining ID and OOD situations:
@@
  • PenInCup (real, 109 demos): The robot, operating on a bare tabletop (no tablecloth), picks up a blue marker pen with its left arm and places it into a red cup. Using a gaze-based action segmentation method, this series of actions is divided into two sub-tasks. The pen is always placed such that the part without the cap remains within a 15 cm square area on the left side of the table, and the cup is placed so that part of its base overlaps a 5 cm area on the right side of the table. The robot’s initial pose starts with its left and right end-effectors positioned above the left and right edges of the table, respectively, oriented inward. The object positions and orientations, as well as the robot’s initial pose, are manually randomized within their respective areas. All other factors (table location, background, etc.) are maintained as consistent as possible during the demonstration collection. To avoid introducing bias in the training images, we marked the boundaries of the object placement areas in a way that would be minimally visible.
@@
  • OpenCap (real, 110 demos): The robot, operating on a tabletop covered with a green tablecloth, holds an upright empty plastic bottle in place with its right arm and removes the cap with its left arm. This series of actions is divided into two sub-tasks. The bottle is always placed within the right side of the table. All other conditions are identical to those described in the PenInCup task.
  • PileBox (sim, 100 demos): The robot, operating on a bare tabletop, picks up a red box with its left arm and stacks it on top of a green box. This series of actions is divided into two sub-tasks. The red box is always placed such that its center remains within a 10 cm × 20 cm rectangular region on the left side of the table, and the green box is placed such that its center remains within a 10 cm square region on the right side of the table. The robot’s initial pose starts with its left and right end-effectors positioned above the left and right edges of the table, respectively, oriented inward. Because this task is performed in a simulation environment, the object positions and orientations are fully randomized according to an uniform distribution, whereas the initial robot pose, table position, and background are all fixed.
IV-C Evaluation of Generalization and Ablation
To verify the improved reusability of learned skills provided by GazeBot, we first trained GazeBot and its ablation models, including conventional models (ACT [29], DAA [15]), using the demonstrations collected as previously described in Section IV-B. We then measured their success rates for scenarios within the training distribution (ID) and those outside it (OOD), as shown in Figure 6. In the experiment, we standardized the initial conditions for all ID and OOD trials to enable an accurate comparison of each model’s generalization performance. Consequently, both GazeBot and the seven ablation models performed the task under conditions as similar as possible. In the OOD trials, we first evaluated the cases where each object was individually placed in unseen positions, and then examined the cases where all objects were simultaneously placed in unseen positions (Figure 6). Finally, we conducted trials involving unseen initial end-effector poses. All initial conditions used in the ID and OOD trials are provided in Appendix -B.
The overall success rates in ID and OOD trials for the PenInCup, OpenCap, and PileBox tasks are presented in Table I. Here, the reusability of acquired skills is evaluated by how well the model achieves a success rate as high as possible on OOD while maintaining a high success rate on ID. GazeBot incorporates several design choices and components, whose contributions can be assessed by comparing its performance with that of the ablation models.
Figure 6: Examples of ID and OOD trials in the PenInCup, OpenCap, and PileBox tasks, where object positions and the initial end-effector poses are controlled. The images show the initial states of each trial. The checkboxes correspond to the method order in Table I and indicate whether each method succeeded in the task from that initial state.
@@ IV-C Evaluation of Generalization and Ablation
GazeBot (Ours) 92 75 83 83 100 83 100 75 100 100 80 75
- w/o point cloud (Ablation1) 75 75 42 42 100 67 58 42 95 65 70 40
- w/ state input partially (Ablation2) 50 33 25 8 100 50 58 42 100 90 85 50
- w/ state input (Ablation3) 75 67 17 0 100 50 50 33 100 80 65 10
- w/o relative bottleneck (Ablation4) 58 50 42 17 100 67 67 42 100 95 65 35
DAA 67 50 17 0 75 67 8 8 100 60 15 0
- w/o local action 67 50 17 0 67 50 8 8 100 40 20 0
ACT 75 67 8 0 100 75 25 25 100 80 40 15
TABLE I: Comparison of success rates (%) for ID and OOD trials in PenInCup, OpenCap and PileBox tasks. Our method is compared with seven ablation models, two of which are conventional baselines. We conducted 12 trials for PenInCup and OpenCap, and 20 trials for PileBox. Notably, only the proposed method maintains a high success rate in OOD situations for both tasks.
@@ IV-C2 Inputs to the Policy (Ablation2, Ablation3)
In GazeBot, the action policy model is designed so that the current end-effector pose is used as input only during the estimation of the bezier vector. In other words, bottleneck estimation relies solely on the 3D gaze position and the gaze-centered point cloud, and the prediction of the gaze-centered dexterous actions after the bottlenecks also relies only on the gaze-centered point cloud.
In Ablation3, we added tokens representing the current left and right end-effector poses and the 3D gaze position (3 × 512-dimensions) to the sequence of 3D point cloud tokens fed to the Transformer encoder. This addition prevented the model from accurately estimating the bottleneck in unseen initial poses. Even when the robot successfully reached the correct bottleneck for unseen object positions, the subsequent gaze-centered dexterous actions degraded because the bottleneck pose itself is an untrained input for the policy network (Table I).
In Ablation2, we similarly added tokens but then applied an attention mask so that the CLS token $f_{pcd}$ in the Transformer encoder did not attend to these additional tokens. As a result, $f_{pcd}$ aggregated only information from the 3D point cloud, enabling accurate bottleneck reaching for unseen end-effector poses. Nonetheless, as mentioned in Ablation3, the accuracy of the gaze-centered dexterous actions still decreased even after the bottleneck was reached correctly (Table I), often causing the end-effector to be “pulled” back toward the ID region.
@@ IV-C4 Design of the Bottleneck Reaching Module (DAA)
In GazeBot, we achieve bottleneck reaching by first predicting a bottleneck pose and then generating a first-order Bézier curve as a trajectory leading to the bottleneck. One could adopt a fully parametric approach that directly outputs an action sequence leading to the bottleneck pose. DAA, for instance, proposes switching between global and local actions—similar to our segmentation by bottlenecks—and employs a fully parametric approach for the global actions by directly predicting the time series of end-effector poses. In our tasks, however, DAA exhibited a significant drop in success rates from ID to OOD (Table I). In most OOD trials, the model even failed to perform the reaching motion itself, reflecting the issues seen in Ablation4.
@@ IV-C5 Gaze-Based Vision (ACT, DAA w/o local action)
Lastly, we compared two baselines: (1) ACT, which is effectively equivalent to removing all proposed modules from GazeBot, and (2) DAA w/o local action, which can be interpreted as applying a gaze-based image-cropping strategy to ACT. As can be observed from Table I, both approaches exhibited a sharp decline in success rates when moving from ID to OOD. In OOD trials, whether local action was present or not had little impact on the success rate of DAA because it already failed during the global (reaching) phase. Moreover, ACT slightly outperformed DAA overall. In DAA, because only a small, gaze-centered region is input to the model and 3D awareness is nearly absent, the input varies significantly with the object position, undermining data efficiency. Consequently, with approximately 100 demonstrations—as in this experiment—it was not possible to achieve the improvements in dexterity previously reported [13]. Meanwhile, GazeBot, which also adopted gaze-based vision, outperformed ACT (Table I). This demonstrates that our approach additionally contributed to improved data efficiency of gaze-based models.
V Limitations and Future Directions
In this study, we proposed GazeBot, an imitation learning method for robotic object manipulation that significantly improves the reusability of learned skills without sacrificing dexterity and reactivity. GazeBot demonstrates a significantly high generalization capability, particularly for unseen object positions and unseen end-effector poses, compared to state-of-the-art models. However, several challenges remain.

### v3

@@ III-C1 Gaze Prediction Model
For the gaze prediction model, a human teleoperator’s gaze is measured simultaneously during demonstration collection, and this gaze data is then used as supervision [7]. The gaze positions estimated by the model are used during both training and inference. As shown in the left side of Figure 4, the image embeddings are extracted via DINOv2 [21], and each 768-dimensional embedding is mapped to a single probability value with a four-layer multilayer perceptron (MLP).
To intentionally transition the gaze to the next gaze landmark upon completing each sub-task, we prepare one such MLP for each sub-task and switch these MLPs based on the sub-task index $i_{\mathrm{seg}}$. At inference time, the sub-task index $i_{\mathrm{seg}}$ is initialized to zero at the beginning and incremented based on the progress $c_{t}$ output by the action policy model, indicating the completion of each sub-task. By managing gaze transitions using $c_{t}$ estimated from the gaze-centered point cloud, which solely captures the relative spatial relationships between the object and end-effector (or grasped object), the robot can perform gaze transitions at the correct timing even under unseen conditions.
@@ III-C2 Action Policy Model (i)
The pixel-space gaze position predicted by the gaze prediction model is transformed into 3D coordinates using stereo-based depth estimation. From the stereo-derived point cloud, we then crop a cubic region (20 cm on each side in this study) centered on the 3D gaze coordinates, which is referred to as the gaze-centered point cloud. This gaze-centered point cloud is embedded as a sequence of tokens using a PointTransformer [22] and then passed to a Transformer encoder (Figure 4 right). The Transformer encoder includes a CLS token (classification token, similar to ViT [23]) that aggregates information from all point-cloud tokens into a feature vector $f_{\mathrm{pcd}}\in\mathbb{R}^{512}$. All output tokens, including the CLS token, are then fed into a Transformer decoder, which generates both the action sequence $a_{t:t+H}\in\mathbb{R}^{H\times 14}$ for the left and right end-effector over $H$ future time steps and the progress $c_{t}\in\mathbb{R}^{N_{seg}}$ for controlling gaze transitions in each sub-task. Here, each end-effector has 7 degrees of freedom including gripper angle, and $N_{seg}$ denotes the number of sub-tasks. As the input vision is gaze-centered, the output actions are represented as relative end-effector poses, with the current end-effector pose serving as the origin of the coordinate frame. By computing these actions in relative coordinates, the learned gaze-centered actions remain reusable even when the absolute position of the object changes.
@@ III-C3 Action Policy Model (ii)
After predicting the bottleneck pose, the corresponding reaching trajectory is generated. Because this trajectory is non-contact and relatively coarse, we model it with a first-order Bézier curve, capturing its overall shape without introducing unnecessary complexity that could hinder generalization. As illustrated in the lower-right portion of Figure 4, the model predicts a 7-dimensional bezier vector that determines the shape of the first-order Bézier curve for each end-effector. A first-order Bézier curve is defined by its two endpoints and a single control point. Rather than predicting the control point directly, we first estimate the displacement—bezier vector—from the mean pose of the start (current end-effector pose) and end (bottleneck pose) poses, and then add this bezier vector to the mean pose to obtain the control point (Figure 4 bottom-right). For the training, we approximate each reaching motion in the demonstrations by fitting them with first-order Bézier curves, and then use the resulting bezier vectors from these curves for supervision.
@@ IV-A Robot System
We used a dual-arm robot system designed for imitation learning via human teleoperation. In this system, a human operator remotely controls the robot while observing its surrounding environment through a head-mounted display (HMD). During this operation, the operator’s gaze data were recorded in sync with the video feed displayed on the HMD. The system is compatible with both a physical dual-arm robot, consisting of two UR5 (Universal Robots Inc.) arms, and its simulated counterpart, allowing the collection of teleoperated demonstration data in both real and virtual environments. The recorded gaze data were output as pixel coordinates corresponding to the images displayed on the HMD. The same single stereo camera (ZED Mini, Stereolabs Inc.) was used for both the teleoperation and inference phases, and we directly utilized the depth images generated by the deep learning–based depth estimation algorithm provided by ZED SDK (Stereolabs Inc.). The time-series demonstration data used for training were recorded at 10 Hz.
Figure 5: Examples of ID and OOD trials in the PenInCup, OpenCap, and PileBox, where object positions and the initial end-effector poses are controlled. The images show the initial states of each trial. The checkboxes correspond to the method order in Table I and indicate whether each method succeeded in the task from that initial state.
@@ IV-B Task Setup
We conducted imitation learning experiments on five tasks (four real and one simulated) to evaluate the skill reusability and applicability of GazeBot.
@@
For the first three tasks, to quantitatively assess the reusability, we collected demonstrations under controlled conditions by restricting the object positions and initial end-effector poses within predefined regions, thus clearly defining ID and OOD situations:
@@
  • PenInCup (real, 109 demos): On a bare tabletop, the robot picks up a blue marker pen with its left arm and places it into a red cup. Gaze-based action segmentation splits this into two sub-tasks. The pen is placed with its uncapped end within a 15 cm square on the table’s left, and the cup base overlaps a 5 cm area on the right. The robot starts with end-effectors above the table edges, oriented inward. Object and robot poses are manually randomized within defined regions; other factors (e.g., table location, background) are kept consistent. Placement boundaries are subtly marked to avoid visual bias.
@@
  • OpenCap (real, 110 demos): On a bare tablecloth, the robot holds an upright plastic bottle with its right arm and unscrews the cap with its left. The task is divided into two sub-tasks. The bottle is always on the right side of the table. Other settings match PenInCup.
  • PileBox (sim, 100 demos): On a bare tabletop, the robot stacks a red box (picked up with its left arm) onto a green box. The task is divided into two sub-tasks. The red box is placed within a 10×20 cm area on the left, and the green box within a 10 cm square on the right. Other settings match PenInCup.
For the remaining two tasks, we chose more complex, long-horizon tasks involving deformable objects to evaluate the model’s applicability (task configurations were approximately randomized):
  • WipeTray (real, 100 demos): On a green tablecloth, the robot folds a towel, regrasp it, and wipes a tray with it, then places the towel back on the table. The task is divided into five sub-tasks.
  • BaggingGoods (real, 91 demos): On a green tablecloth, the robot grasps the handle of a bag with its left arm and opens the bag. It then uses its right arm to place two of the three objects on the table into the bag in a specified order. The task is divided into five sub-tasks.
IV-C Evaluation of Reusability and Ablation Studies
TABLE I: Comparison of success rates for ID and OOD trials in PenInCup, OpenCap and PileBox.
@@ IV-C Evaluation of Reusability and Ablation Studies
GazeBot (Ours) 11/12 9/12 10/12 10/12 12/12 10/12 12/12 9/12 20/20 20/20 16/20 15/20
- w/o point cloud (Ablation1) 9/12 9/12 5/12 5/12 12/12 8/12 7/12 5/12 19/20 13/20 14/20 8/20
- w/ state input partially (Ablation2) 6/12 4/12 3/12 1/12 12/12 6/12 7/12 5/12 20/20 18/20 17/20 10/20
- w/ state input (Ablation3) 9/12 8/12 2/12 0/12 12/12 6/12 6/12 4/12 20/20 16/20 13/20 2/20
- w/o relative bottleneck (Ablation4) 7/12 6/12 5/12 2/12 12/12 8/12 8/12 5/12 20/20 19/20 13/20 7/20
DAA 8/12 6/12 2/12 0/12 9/12 8/12 1/12 1/12 20/20 12/20 3/20 0/20
- w/o local action 8/12 6/12 2/12 0/12 8/12 6/12 1/12 1/12 20/20 8/20 4/20 0/20
ACT 9/12 8/12 1/12 0/12 12/12 9/12 3/12 3/12 20/20 16/20 8/20 3/20
We first trained GazeBot and its ablation models, including conventional models (ACT [1], DAA [4]), using the demonstrations collected as previously described. ACT is known for its strong performance and can be considered equivalent to GazeBot with all proposed modules removed. DAA extends ACT by incorporating gaze-based image cropping and action segmentation into global and local actions. DAA w/o local actions can be interpreted as applying only gaze-based vision to ACT. We then measured their success rates for scenarios within the training distribution (ID) and those outside it (OOD), as shown in Figure 5. In the experiment, we standardized the initial conditions across all trials to ensure that all models performed the task under as similar conditions as possible. In the OOD trials, we first evaluated the cases where each object was individually placed in unseen positions, and then examined the cases where all objects were simultaneously placed in unseen positions (Figure 5). Finally, we conducted trials involving unseen initial end-effector poses.
The overall success rates in ID and OOD trials for the four tasks are presented in Table I. Here, the reusability of acquired skills is evaluated by how well the model achieves a success rate as high as possible on OOD while maintaining a high success rate on ID. GazeBot incorporates several design choices and components, whose contributions can be assessed by comparing its performance with that of the ablation models.
@@ IV-C2 Inputs to the Policy (Ablation2, Ablation3)
In GazeBot, the action policy model is designed so that the current end-effector pose is used as input only during the estimation of the bezier vector. In other words, bottleneck estimation relies solely on the 3D gaze position and the gaze-centered point cloud, and the prediction of the gaze-centered actions after the bottlenecks also relies only on the gaze-centered point cloud.
In Ablation3, we added tokens representing the current left and right end-effector poses and the 3D gaze position (3 × 512-dimensions) to the sequence of 3D point cloud tokens fed to the Transformer encoder. This addition prevented the model from accurately estimating the bottleneck in unseen initial poses. Even when the robot successfully reached the correct bottleneck for unseen object positions, the subsequent gaze-centered actions degraded because the bottleneck pose itself is an untrained input for the policy network (Table I).
In Ablation2, we similarly added tokens but then applied an attention mask so that the CLS token $f_{pcd}$ in the Transformer encoder did not attend to these additional tokens. As a result, $f_{pcd}$ aggregated only information from the 3D point cloud, enabling accurate bottleneck reaching for unseen end-effector poses. Nonetheless, as mentioned in Ablation3, the accuracy of the gaze-centered actions still decreased even after the bottleneck was reached correctly (Table I), often causing the end-effector to be “pulled” back toward the ID region.
@@ IV-C4 Design of the Bottleneck Reaching Module (DAA)
In GazeBot, we achieve bottleneck reaching by first predicting a bottleneck pose and then generating a first-order Bézier curve as a trajectory leading to the bottleneck. One could adopt a fully parametric approach that directly outputs an action sequence leading to the bottleneck pose, as in DAA. In our tasks, however, DAA exhibited a significant drop in success rates from ID to OOD (Table I). In most OOD trials, the model even failed to perform the reaching motion itself, reflecting the issues seen in Ablation4.
@@ IV-C5 Gaze-Based Vision (ACT, DAA w/o local action)
As can be observed from Table I, both ACT and DAA exhibited a sharp decline in success rates when moving from ID to OOD. In OOD trials, whether local action was present or not had little impact on the success rate of DAA because it already failed during the global (reaching) phase. Moreover, ACT slightly outperformed DAA overall. In DAA, because only a small, gaze-centered region is input to the model and 3D awareness is nearly absent, the input varies significantly with the object position, undermining data efficiency. Consequently, with approximately 100 demonstrations, as in this experiment, it was not possible to achieve the improvements in dexterity previously reported [8]. Meanwhile, GazeBot, which also adopted gaze-based vision, outperformed ACT (Table I). This demonstrates that our approach additionally contributed to the improved data efficiency of gaze-based models.
TABLE II: Comparison of success rates in WipeTray and BaggingGoods.
WipeTray (real) BaggingGoods (real)
Method Folded Wipe Open Insert1 Insert2
GazeBot (Ours) 17/20 16/20 17/20 12/20 11/20
ACT 18/20 18/20 15/20 7/20 7/20
V Evaluation of Applicability
We evaluated skill reusability using simple rigid-body tasks to establish a clear ID/OOD split. However, the aim of our study is to enhance reusability without limiting the task scope to basic pick-and-place operations involving only rigid objects. Therefore, in this section, we demonstrate that GazeBot can also successfully learn more complex and realistic tasks, WipeTray and BaggingGoods, both of which involve long-horizon and deformable-object manipulation. Tests were conducted under demonstration-like configurations without defining ID/OOD. As a baseline, we used ACT, one of the most capable model-free approaches. In each trial, we ensured that all models operated under the same configuration as much as possible.
The results are summarized in Table II. GazeBot successfully learned all tasks, similar to ACT. Moreover, it clearly outperformed ACT in BaggingGoods, likely because this task involves many objects and large variations in their configurations. In such cases, GazeBot’s gaze- and bottleneck-based canonicalization helps improve learning accuracy. In contrast, ACT performed better in WipeTray, where the object is large and its position varies little. In these scenarios, ACT can more easily reproduce the demonstrated behavior by directly incorporating proprioception into the action policy.
VI Limitations and Future Directions
In this study, we proposed GazeBot, an imitation learning method that significantly improves skill reusability without sacrificing dexterity or reactivity, particularly for unseen object positions and end-effector poses. However, several challenges remain.